OpenStudy (anonymous):

A baker is caculating the charge for two types of cookies what formula tells the cost in dollars if chocolate chip cookies are $ 2.50 / dozen and lemon frosted are $ 1.50 / dozen? Let c = number of dozens of chocolate chip cookies L = Number of dozens of lemon frosted t = total charge

OpenStudy (anonymous):

It's multiple choice so these are the only options I have to choose from, T=250c + 150L, T=400 (L + c, T=1.50 + 2.50L, T = 2.50c + 1.50L

OpenStudy (anonymous):

T=4.00 (L +c

OpenStudy (anonymous):

2.5L+1.5C=T

OpenStudy (anonymous):

i think you made a mistake when you wrote the 3rd choice..

OpenStudy (anonymous):

Yeah lol